I need help on how to turn this word problem into an equation. "A theatre has 25 seats in the first row and 35 rows in all. Each successive row contains one additional seat. How many seats are in the theater?"
Answer by stanbon(75887)

"A theatre has 25 seats in the first row and 35 rows in all. Each successive row contains one additional seat. How many seats are in the theater?"
---------------------------
Arithmetic Series with a(1) = 25 and d = 1
--------
S(n) = a(1) + (n-1)d
-----
S(35) = 25 + 34*1 = 59 seats
======================
Cheers,
Stan H.
======================